iven
f(x) = 2x2 + 1
and
g(x) = 5x − 9,
find the following.
(a)
f(g(2))
(d)
(g ∘ g)(x)

Answer by rothauserc(4718)

(f o g)(x) = 2(5x-9)^2 +1 = 2(25x^2 -90x +81) +1 = 50x^2 -180x +162 +1 =
50x^2 -180x +163
:
(a) f(g(2) = 50 * 2^2 -180 * 2 +163 = 3
:
(d) (g o g)(x) = 5(5x-9) -9 = 25x -45 -9 = 25x -54
: